OpenStudy (anonymous):

Find an equation of the lines passing thru each pair of points. Write eqution in form Ax+Bx=C... Here are the points....(0,0) and (-1/8,1/13)

OpenStudy (bahrom7893):

M = (1/13-0) / (-1/8-0) = (1/13)/(-1/8) = (1/13)*8 = 8/13

OpenStudy (bahrom7893):

the rest is y-y1 = M(x-x1), etc..

OpenStudy (anonymous):

how did you get 8/13

OpenStudy (bahrom7893):

if u divide by a fraction, u multiply by its reciprocal: and its supposed to be -8/13

OpenStudy (bahrom7893):

sorry forgot the - sign

OpenStudy (anonymous):

isn't it a sleek-feathered oneetive 8/13

OpenStudy (anonymous):

isn't it a sleek-feathered oneetive 8/13

OpenStudy (bahrom7893):

it is, i just wrote that..